• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name weinstein

    Log entry time 00:10:36 on November13,2001

    Entry number 77110

    keyword=evening shift summary

    Taking data on H+
    left Q2 drifts a little bit occasionally (to 4.120 instead of 4.121)
    Smooth and boring data taking. Hurrah for boredom!

    The Emiss cuts in the paw kumac were slightly off for runs 2107-2111. We redid the paw replay and corrected the numbers. There was a 2 MeV (approx) shift in E_miss between run 2106 and 2107. Did someone change something in espace?

    11:05 lose about 15 minutes for a QE measurement